1) Determine the oxidation number if sulfur in the SO⅔-ion and the SO²4 -ion and the use number to name Each ion.


1
Expert's answer
2022-01-21T02:37:05-0500

SO32− ion: S+4; O−2

1×(+4) + 3×(−2) = 4 + (−6) = (−2) – charge of ion (correct)

The oxidation number of S in SO32− ion is +4

SO32− is the sulfite ion


SO42− ion: S+6; O−2

1×(+6) + 4×(−2) = 6 + (−8) = (−2) – charge of ion (correct)

The oxidation number of S in SO42− ion is +6

SO42− is the sulfate ion